a. John won a lottery worth $145,000 and she is expecting to buy a business in five years from now. He has been told by the bank that he would need about $350,000 to set the business in the future. What percentage of the business can be funded by the money from the lottery in 5 years from now. [Assume that interest will remain 7.7% unchanged. ]

b. Mr LEO’s father has made a will to give him an amount of $500,000 in 10 years .So Leo wants to borrow an amount from the bank to set up a fashion boutique in the nearest town. The business planner told him that he does not expect the present discount rate to change in these 10 years. The following information is given regarding the discount rate

  GDP growth of the country

1.2%

Beta

1.8

Index in January 2020

16,570

Index in December 2020

17,590

Inflation

2.5%

 REQUIRED

How much should he borrow from the bank now if he is planning to pay back the loan with the money he will get in 10 years of business return?


1
Expert's answer
2021-06-07T17:31:09-0400

,A) percent of the business in 5 years=

=(P-I)====> ($350000-$145000)=$205,000

(A×R/100×T)===>($205,000×7.7/100×5)= $14350

B) (A×T) where the previous will be divided by the second year that is 10/5=2

= ($14350×2)= $28,700, the given information in part B question only explain the list of items items purchased and is not Applicable in solving the questions